Find the expansion of


a) (x + y)^6


b) (x + y)^4


31) What is the coefficient of x^12 y^13 in the expansion of (x + y)^25?


32) What is the coefficient of x9 in (2 − x)19?


33) What is the coefficient of x^101 y^99 in the expansion of (2x − 3y)^200?

1
Expert's answer
2021-12-24T14:19:10-0500

a)


"(x+y)^6=x^6+6x^5y+15x^4y^2+20x^3y^3+15x^2y^4"

"+6xy^5+y^6"

b)


"(x+y)^4=x^4+4x^3y+6x^2y^2+4xy^3+y^4"

31)


"\dbinom{25}{12}=\dfrac{25!}{12!(25-12)!}=5200300"

32)


"(-1)^9\dbinom{19}{10}(2)^{10}=-\dfrac{19!}{10!(19-10)!}(1024)"

"=-92378(1024)=-94595072"

33)


"(-1)^{99}\dbinom{200}{101}(2)^{101}(3)^{99}"

"=-\dfrac{200!}{101!(200-101)!}(2)^{101}(3)^{99}"